Vector a is 6m due north, b is 4 at an angle of 30 degree Celsius above the horizontal what is the resultant vector?


1
Expert's answer
2022-01-18T09:35:13-0500
"R_N=6+4\sin30°=8\text{ m},\\
R_E=0+4\cos30°=3.46\text{ m},\\
R=\sqrt{R_N^2+R_E^2}=8.72\text{ m},\\\space\\
\theta=\arctan\dfrac{R_N}{R_E}=66.6°."